OpenStudy (anonymous):

Suppose p varies directly as d, and p = 2 when d = 7. What is the value of d when p = 10? A.)d=35 B.)d=15 C.)d=20/7 D.)d=7/5

OpenStudy (diyadiya):

p varies directly with d so this equation can be written as:- p=\(k\)d where \(k\) is the constant of variation 2=7\(k\) \(k\)=2/7 so when p is 10 10=\(k\)d 10=2/7*d d=35